two impedances z1=3+j6 ohms and z2=4-j3 ohms are connected in series. what is the magnitude of the resultant impedance ?
Answer #1
Two impedances Z1=4+j3 and Z2= 3+j4 are connected in parallel. The resultant impedance is____